OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 10 = x2 + 3x?

OpenStudy (luigi0210):

the x-intercept is when y=0

OpenStudy (luigi0210):

Uh, no..

OpenStudy (anonymous):

10=x2+3x is were i am at

OpenStudy (luigi0210):

Correct, now now that 10 over, set it equal to 0 and factor

OpenStudy (anonymous):

so 0=x2 +3x?

OpenStudy (luigi0210):

Where's your 10? o:

OpenStudy (anonymous):

you said make it equal to 0

OpenStudy (luigi0210):

Yes, but that doesn't mean make your 10 disappear into thin air xD

OpenStudy (anonymous):

oh were does it go?

OpenStudy (luigi0210):

To the other side, you subtract it from both sides

OpenStudy (anonymous):

0=x2+3x-10

OpenStudy (luigi0210):

Corrrect, now factor

OpenStudy (anonymous):

nothing goes into all of them

OpenStudy (anonymous):

wait

OpenStudy (luigi0210):

Yes there is

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

x=2, x=-5?

OpenStudy (luigi0210):

Yea, good job :)

OpenStudy (anonymous):

ty

OpenStudy (luigi0210):

btw Welcome to openstudy :) If you have any questions concerning the site feel free to PM me or any other ambie with an A next to their name. And just for future references, if someone helps you out you can give them a medal by hitting the "Best Response" button next to their name. Just saying because there are some users on here who get annoyed when they aren't awarded or something :P Anyways enjoy the site! :D

OpenStudy (anonymous):

whats an ambie?

OpenStudy (luigi0210):

Ambie is short for Ambassador, and out job is to help and welcome new users to the site. Making them feel comfortable and welcomed :)

OpenStudy (luigi0210):

*our

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

well ty for the help

OpenStudy (luigi0210):

You are very welcome

OpenStudy (anonymous):

can you help me with one quick one?

OpenStudy (anonymous):

What are the solutions of 5x2 = -15x?

OpenStudy (anonymous):

nvm i got it 5 , -3

OpenStudy (luigi0210):

Actually there should only be one solution

OpenStudy (luigi0210):

One of the answers should be 0